OpenStudy (anonymous):

In order to do pathygreom therom how do i know which is A,B,C of the triangle...?

OpenStudy (amistre64):

the diagram tells you; or the problem tells you

OpenStudy (amistre64):

dont get hung up on the 'names' of the variables; just know what the pythag meas

OpenStudy (anonymous):

The shortest side is A, the side perpendicular to that is B, and the diagonal is C.

OpenStudy (amistre64):

leg^2 + leg^2 = hyp^2

OpenStudy (anonymous):

Here is the problem... maybe this whill help out with my question a little

OpenStudy (amistre64):

this is the application of the law of cosines

OpenStudy (anonymous):

But dont i need to find out the other side in order to do law of cosines..?

OpenStudy (anonymous):

ohh nevermind your right!!!! :]] thankss

OpenStudy (amistre64):

the third side is: 3rd side = \(\sqrt{6^2+10^2 -(2*6*10*cos(80))}\)

OpenStudy (amistre64):

once you have that sides; you can law of sine it to ge the angles

OpenStudy (anonymous):

thats the c^2=a^2*b^2-2(a)(c)cos80 yea :] but couldnt i use pathygreom therom for this ...

OpenStudy (amistre64):

well, that IS the pythag thrm, or better; the pythag thrm is the law of cosines when the angle = 90

OpenStudy (amistre64):

what happen when cos(90) is used in the law of cosines?

OpenStudy (amistre64):

since you only know one angle, and it aint 90; pythag aint gonna help you out is it

OpenStudy (anonymous):

ohh well i guess that explains it alot i just use it all the time no matter what the circumstance is if i only have two angles i usually just do pathagreom theroem though im always wrong i never know why youve been much help as always to me thank you !

OpenStudy (amistre64):

:) youre welcome